Two particles 1 and 2 are thrown in the direction downward and upward with velocity 5m/s and 20m/s. Initially particle 1 is at height 20m from ground. Taking upward direction as positive direction.


Find (a)acceleration of 1 w.r.t 2


(b)Initial velocity of 2 w.r.t 1


(c)velocity of 1 w.r.t 2 after t=0.5s


(d)time when the particle will collide?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

a) relative accleration of 1 wrt 2 is A12=g-(-g)=2g

b)velocity of 2 wrt 1 is U21=u2-u1=20-(-5)=25m/s

c)after 0.5 sec velocity of first is v1=-5-5=-10m/s

   velocity of second is v2=20-5=15m/s

V12=V1-V2=-10-(15)=-25m/s

d)let distance travelled by particle in time t is -S1=-5t-5t^2

                    S1=5t+5t^2.....1

let distance travelled by particle 2 after time t is S2=20t-5t^2.......2

S1+S2=20           

25t=20

t=4/5 sec

at t=4/5 sec particle will collide